OpenStudy (anonymous):

Completely factor the expression 147x^2y+42xy^2+3y^3

hero (hero):

Wow, what a problem

hero (hero):

3y(49x^2+14xy+y^2) = 3y(49x^2 +7xy + 7xy +y^2) = 3y(7x(7x+y) + y(7x + y)) = 3y((7x+y)(7x+y)) = 3y(7x+y)^2